the midpoint of a segment is (2, -5), and one of the endpoints is (3, 6). Where is the other endpoint?
Answer by jim_thompson5910(35256)

(x+3)/2 = 2


x+3 = 4


x = 1

-------------------------------------------------------

(y+6)/2 = -5


y+6 = -10


y = -16



So the other endpoint is (1, -16)